LINE INTENSITIES MEASUREMENTS IN ^{13}C^{16}O_{2} AND THEIR TREATMENT USING THE EFFECTIVE OPERATORS APPROACH

V. I. PEREVALOV, S. A. TASHKUN, Institute of Atmospheric Optics, Russian Academy of Sciences, Siberian Branch, 1 Akademicheskii Avenue, 634055 Tomsk, Russia; C. CLAVEAU, J–L. TEFFO, Laboratoire de Physique Moléculaire pour l’Atmosphère et l’Astrophysique, CNRS UMR 7092, case 76, Université Pierre et Marie Curie, 4 Place Jussieu, F-75252 Paris Cedex 05, France; \underline{J. VANDER AUWERA}, Service de Chimie Quantique et Photophysique C.P. 160/09, Université Libre de Bruxelles, 50 Avenue F. D. Roosevelt, B–1050 Brussels, Belgium.

This work deals with the modeling of line intensities of cold and hot bands of ^{13}C^{16}O_{2} corresponding to ΔP = 5 transitions (P = 2V_{1} + V_{2} + 3V_{3}), using the effective operators approach.^{a} In Brussels, 5 unapodized absorption spectra of ^{13}C^{16}O_{2} (98% purity, P × ℓ = 7.43 − 538 mbar×m, room temperature) were recorded at a resolution of 0.0035 cm^{−1} (MOPD = 257.1 cm) using a Bruker IFS120HR. 382 absolute line intensities were measured in 9 bands observed between 3090 and 3920 cm^{−1}. In Paris, 2 apodized absorption spectra of an ^{18}O–enriched CO_{2} sample (297.0 K, ℓ = 8.07(2) and 36.07(4) m, P = 1.685(8) torr) containing 0.5% of ^{13}C^{16}O_{2} were recorded at a resolution of 0.0043 cm^{−1} using a home– made FTS. 90 absolute line intensities were measured for 5 bands observed from 3400 to 3700 cm^{−1}. A least squares fit of 479 observed line intensities, including those measured in this work and published for the 21102-00001 band,^{b} allowed determination of 12 effective dipole moment parameters corresponding to the ΔP = 5 series of transitions (weighted standard deviation χ = 0.855, RMS deviation = 4.69%). The new set of effective dipole moment parameters will soon be used to update the Carbon Dioxide Spectroscopic Databank (CDSD–1000).^{c} Results will be presented and discussed.

aS.A. Tashkun, V.I. Perevalov, J.-L. Teffo, L.S. Rothman, and Vl.G. Tyuterev, J. Quant. Spectrosc. Radiat. Transfer \underline{60}, 785–801 (1998); J.-L. Teffo, O.M. Lyulin, V.I. Perevalov, and E.I. Lobodenko, J. Mol. Spectrosc. \underline{187}, 28–41 (1998).

bY. Ding, P. Macko, D. Romanini, V.I. Perevalov, S.A. Tashkun, J.-L. Teffo, S. Hu, and A. Campargue, J. Mol. Spectrosc. \underline{226}, 146–160 (2004).

cS.A. Tashkun, V.I. Perevalov, J-L. Teffo, A.D. Bykov, and N.N. Lavrentieva, J. Quant. Spectrosc. Radiat. Transfer \underline{82}, 165–196 (2003); http://cdsd.iao.ru, http://cdsd.lpma.jussieu.fr, ftp://ftp.iao.ru/pub/CDSD-1000.